# Nonconvergence of the Feynman–Dyson diagrammatic perturbation expansion of propagators

### v7

@@ IV.1 Formalisms
To avoid malaise of truncated perturbation approximations, infinite partial summations of diagrams have been tried in the past. One way of doing this is by renormalizing vertexes. (Here, a “vertex” is an instantaneous Coulomb interaction denoted by a horizontal line rather than a spacetime point of contact between propagator lines and an interaction line.) In the two-particle-hole Tamm–Dancoff approximation (TDA) Linderberg and Öhrn (1967); Purvis and Öhrn (1974); Schirmer and Cederbaum (1978); Walter and Schirmer (1981), also known as the Brueckner–Hartree–Fock method Day (1967); Baldo (1999) or the T approximation Kadanoff and Baym (2018), the diagrams of the types in Fig. 16 are summed over up to an infinite order. In the figure, the first two diagrams in the third and fourth lines are “ring” diagrams, and the subsequent two diagrams are “ladder” diagrams. The effect of this infinite summation can be folded into the bold-line vertexes as appearing in the first line of Fig. 16, which must, in turn, satisfy the diagrammatic equations drawn in Fig. 17. They take the form of the amplitude equations of (linearized) coupled-cluster theory Bartlett and Musiał (2007); Shavitt and Bartlett (2009) and are written algebraically as
@@
Equations (28) and (29) are a system of linear equations, which can, therefore, be ill-conditioned at some $\omega$’s. In this work, they are solved in an iterative algorithm; that is, starting with initial guesses of $\bm{U}$ and $\bm{V}$, we substitute them in the right-hand sides of these equations to update $\bm{U}$ and $\bm{V}$ in the left-hand sides, and repeat this process until convergence. Therefore, in practice, the highest order of the ladder and ring diagrams that are actually included in the calculation is capped by the number of cycles taken in this iterative solution. Upon convergence, the self-energy is obtained as
@@
See Ref. Opoku et al. (2023b) for a more efficient and stable algorithm, which, however, is not exempt from the pathological behaviors discussed below, which are fundamental.
@@
For a total energy, an infinite partial summation of essentially the same ladder and ring diagrams defines an instance of coupled-cluster theory Bartlett and Musiał (2007); Shavitt and Bartlett (2009) known as linearized coupled-cluster doubles (LCCD) or by other names Čížek (1966); Hurley (1976); Bartlett and Shavitt (1977, 1978); Bartlett et al. (1979); Ahlrichs (1979); Koch and Kutzelnigg (1981). This method is $\omega$-independent and is thus free from the nonconvergence problems discussed here. It instead displays other problems concerning accuracy Taube and Bartlett (2009).
Figure 17: The diagrammatic equations for the bold-line vertexes of TDA.
@@
Another way of performing an infinite partial summation of diagrams is by renormalizing edges. By replacing all three edges in each diagram of $\delta\bm{\Sigma}^{(2)}$ by the corresponding bold-line Green’s functions $\bm{G}^{\text{Dyson}(2)}$ of Fig. 7, we include an infinite number of “row-house” diagrams appearing in the second line of Fig. 18. Furthermore, if the bold-line Green’s function (designated by $\bm{G}^{\text{sc}(2)}$) replacing the edges of the self-energy comes from this very edge-modified self-energy itself $\bm{\Sigma}^{\text{sc}(2)}$, as in Fig. 19, we then account for another infinite set of “tower” diagrams shown in the third line of Fig. 18. This self-consistency between the self-energy and Green’s function was emphasized by Baym and Kadanoff Baym and Kadanoff (1961); Baym (1962); Kadanoff and Baym (2018); Dahlen and van Leeuwen (2005) as an essential ingredient of an approximate MBGF method that obeys conservation laws. (The first-order self-consistent MBGF method is identified as the HF theory, which obeys conservation laws Dickhoff (1999).) In chemistry, approximations inspired by this idea are enjoying a revival for its possible ability to describe strong correlation more accurately Surján and Szabados (1998); Holleboom and Snijders (1990); Phillips and Zgid (2014); Coveney and Tew (2023).
@@
Figure 19: The self-consistent second-order bold-line Green’s function.
@@
This ansatz may differ somewhat from those of Van Neck et al. Van Neck et al. (1991) or of Dahlen and van Leeuwen Dahlen and van Leeuwen (2005) as it involves some additional approximations, but they have no impact on the following analysis. The corresponding residues are given by
@@
IV.2 Numerical results
Figure 20: The exact ${\epsilon_{3}}+{\Sigma_{33}}(\omega)$ and ${\epsilon_{3}}+\Sigma_{33}^{\text{TDA}}(\omega)$ as functions of $\omega$ for the BH molecule. The self-energy in TDA is determined after $n$ cycles of the iterative solution of the amplitude equation, summing over the ladder and ring diagrams through the order $n+1$. The 2p1h and 2h1p HF orbital energy differences are superposed.
Figure 20 compares the self-energy of TDA with the exact self-energy. The former was obtained after either 20 or 21 cycles of the iterative solution of the amplitude equations, summing over ladder and ring diagrams through the order 21 or 22, respectively. As may be expected from the discussion on perturbative self-energies, TDA does not improve the overall appearance of the self-energy; rather, it seems to degrade it.
For instance, $\bm{\Sigma}^{\text{TDA}}$ obtained after 20 cycles has qualitatively wrong (concave and convex) functional forms (outside the central bracket) because computationally it is a high-odd-order perturbation theory. In contrast, $\bm{\Sigma}^{\text{TDA}}$ after 21 cycles has qualitatively correct functional forms (for it is a high-even-order perturbation theory), but they consist of near vertical (twenty-first-order) singularities at the 2p1h and 2h1p HF orbital energy differences (which are superposed in Fig. 20). Therefore, TDA predicts vastly different IPs and EAs depending on the number of cycles taken in the iterative solution—an artifact of calculations—and is therefore methodologically ill-defined insofar as all roots are concerned. Furthermore, both of these conflicting predictions are equally meaningless. After 20 cycles, there are no real roots outside the central bracket in this graph; after 21 cycles, roots outside the central bracket tend to coincide with the sheer 2p1h or 2h1p HF orbital energy differences with null correlation effects.
While it is unsurprising for an infinite diagram summation to exacerbate the pathologies of higher-order perturbation theory, it may be too hasty to conclude that the same pathologies plague all diagram summations. The equation-of-motion (EOM) formalism of MBGF Rowe (1968); Simons and Smith (1973); Herman et al. (1978, 1980a, 1980b) also involves an infinite partial summation of diagrams, but it can be recast into a matrix diagonalization Baker and Pickup (1980) and may thus be free from the type of the nonconvergence problems discussed above. The random-phase approximation (RPA) Dickhoff (1999), which is a lowest-order member of the EOM approximation series, as well as the related GW Hedin (1965); Hybertsen and Louie (1986); Aryasetiawan and Gunnarsson (1998); Onida et al. (2002) and frequency-dependent part of the ADC self-energy Schirmer (1982); Schirmer et al. (1983) may also be more robust than TDA, although this has not been numerically confirmed here. The Parquet method Bergli and Hjorth-Jensen (2011) sums over a related, but distinct class of diagrams and hence the present conclusion for TDA does not immediately apply.
@@ IV.2 Numerical results
Figure 21 compares the self-energies of the self-consistent second-order Green’s-function method after zeroth, first, and second iterative cycles. In each case, the self-energy has the qualitatively correct functional form, that is, monotonically decreasing in each bracket separated by its singularities. This is because the self-energy expression [Eq. (32)] is isomorphic to the second-order self-energy, which has only the first-order poles and thus a qualitatively correct form. Furthermore, unlike higher-order perturbative self-energies or TDA, the brackets’ boundaries shift from one cycle to the next, and hence the satellite roots no longer have to approach mere HF orbital energy differences with no correlation.
@@
Figure 22 shows the poles and residues in the zeroth and first cycles of the self-consistent second-order Green’s-function method in comparison with the exact poles and residues. The bottom panel is the same as Fig. 4, but shown in a full domain of $\omega$, and the top panel is equivalent to MBGF(2) in the diagonal approximation. In the zeroth cycle ($n=0$), the distribution of the poles agree reasonably well with that of the exact poles. The number of poles is 72 as compared with 600 exact poles including ones with zero residues. In the first cycle ($n=1$), the number of poles already reaches 4,314, far exceeding the total number (600) of ionized and electron-attached states of a FCI calculation. Furthermore, these spurious poles are not necessarily “phantom” mathematical roots with zero residues; they have nonzero residues and encroach on the regions where there are no exact poles. In the second cycle ($n=2$), the number of roots reaches such an astronomical value that our computer code can no longer handle, and we judged that it was not worthwhile to pursue full self-consistency. This difficulty was recognized by earlier workers, many of whom then decided to abandon this class of methods.
The bold-line self-energy diagrams in Fig. 18 is obtained by cutting one bold line (and then trimming the resulting two dangling lines) Hirata et al. (2017) of a skeleton diagram of the Luttinger–Ward functional Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Gunnarsson et al. (2017); Lin and Lindsey (2018). Several mathematical difficulties associated with this functional have been discovered recently Schäfer et al. (2013); Stan et al. (2015); Kozik et al. (2015); Rossi and Werner (2015); Schäfer et al. (2016); Tarantino et al. (2017); Gunnarsson et al. (2017).
V Discussion
What is the root cause of the nonconvergence? This question can be answered by analyzing a model Green’s function of the form Hirata (2021b),
@@ V Discussion
which consists of four poles. These poles, in turn, depend on the perturbation $\lambda$ as
@@
The functional forms of $g(\omega)$ and $E$’s are arbitrary, but a different choice would lead to the same conclusion.
@@
A truncation of this series after a finite number of terms captures essential features of the Feynman–Dyson perturbation approximations of a Green’s function or self-energy.
Figure 23: Taylor expansions up to the second order in $\lambda$ of Eq. (37).
Figure 23 shows the zeroth-, first-, and second-order Taylor expansions of $g(\omega)$. The exact $g(\omega)$ and its zeroth-order approximation $g^{(0)}(\omega)$ have essentially the same functional forms in that they are separated into consecutive $\omega$ brackets by their singularities and within each bracket they are monotonically decreasing functions. The first-order approximation $g^{(0)}+g^{(1)}$ has a qualitatively different functional form, which is convex or concave except in the central bracket. The $g^{(0)}+g^{(1)}+g^{(2)}$ largely restores the same functional form as the exact $g(\omega)$, although the former is a rather poor approximation at many frequencies despite the fact that the second-order Taylor expansions of $E_{1}$ through $E_{4}$ are exact. Nonetheless, these are consistent with the overall patterns of behaviors of ab initio odd- and even-order perturbative self-energies observed in the previous sections.
@@
Figure 24 extends this analysis to the nineteenth-order Taylor expansion. It is convergent at the exact $g(\omega)$ in some domains of $\omega$, but nonconvergent in the other domains. Generally, convergence is attained within the overlap of the bracket demarcated by the singularities of $g(\omega)$ and the one delineated by the singularities of $g^{(0)}(\omega)$. (The latter singularities are where the near-vertical rapid oscillations of the nineteenth-order Taylor expansion occur in this figure.) In this overlap, both $g(\omega)$ and $g^{(0)}(\omega)$ are in the middle “shoulder” part of the monotonically decreasing functions of $\omega$, displaying similar functional forms. Here, a Taylor expansion from $g^{(0)}(\omega)$ is convergent. Outside this overlap, either $g(\omega)$ or $g^{(0)}(\omega)$ is in the left “neck” part of the function, while the other is in the right “arm” part. They are both monotonically decreasing, but one of them has a singularity at one boundary of the domain, while the other has a singularity at the other boundary, and they have dissimilar functional forms in this regard. The $g(\omega)$ in this domain is nonanalytic and its Taylor expansion has zero radius of convergence.
@@
In a molecular Green’s function or self-energy, there are dense manifolds of singularities outside the central overlapping bracket, which encloses most principal roots and some low-energy satellite roots. Therefore, in practice, their perturbation expansions are safely convergent only in this central overlapping bracket as well as in the two terminal overlapping brackets, although this does not preclude other overlapping brackets in which the perturbation series is convergent. The central overlapping bracket is an overlap of the bracket bounded by the highest 2p1h and lowest 2h1p HF orbital energy differences and the bracket demarcated typically by the least negative and least positive singularities of the exact self-energy Sun and Bartlett (1996b). Only the roots in this central overlapping bracket, be they principal or satellite, are reliably determined by MBGF.

### v8

@@ IV.1 Vertex renormalization
In the two-particle-hole Tamm–Dancoff approximation (TDA) Linderberg and Öhrn (1967); Purvis and Öhrn (1974); Schirmer and Cederbaum (1978); Walter and Schirmer (1981), also known as the Brueckner–Hartree–Fock method Day (1967); Baldo (1999) or the T approximation Kadanoff and Baym (2018), the diagrams of the types in Fig. 16 are summed over up to an infinite order. In this figure, the first two diagrams in the third and fourth lines are “ring” diagrams, and the subsequent two diagrams are “ladder” diagrams. The effect of this infinite resummation can be folded into the bold-line vertexes as appearing in the first line of Fig. 16, which must, in turn, satisfy the diagrammatic equations drawn in Fig. 17. They take the form of the amplitude equations of coupled-cluster theory Bartlett and Musiał (2007); Shavitt and Bartlett (2009) and are written algebraically as
@@
Equations (28) and (29) are a system of linear equations, which can, therefore, be ill-conditioned at some $\omega$’s. In this work, they are solved in an iterative algorithm; i.e., starting with initial guesses of $\bm{U}$ and $\bm{V}$, we substitute them in the right-hand sides of these equations to update $\bm{U}$ and $\bm{V}$ in the left-hand sides, and repeat this process until convergence. Therefore, in practice, the highest order of the ring and ladder diagrams that are actually included in the calculation is capped by the number of cycles taken in this iterative solution. Upon convergence, the self-energy is obtained as
@@
See Ref. Opoku et al. (2023b) for a more efficient and stable algorithm, which, however, is not exempt from the pathological behaviors discussed below, which are deeply rooted in the formalism.
@@
For a total energy, an infinite partial resummation of the corresponding ring and ladder diagrams defines an instance of coupled-cluster theory Bartlett and Musiał (2007); Shavitt and Bartlett (2009) known as linearized coupled-cluster doubles (LCCD) or by other names Čížek (1966); Hurley (1976); Bartlett and Shavitt (1977, 1978); Bartlett et al. (1979); Ahlrichs (1979); Koch and Kutzelnigg (1981). This method is $\omega$-independent and is thus free from the nonconvergence problems discussed here. It instead displays other problems concerning accuracy Taube and Bartlett (2009).
Figure 17: The diagrammatic equations for the bold-line (renormalized) vertexes of TDA.
Figure 18: The exact ${\epsilon_{3}}+{\Sigma_{33}}(\omega)$ and ${\epsilon_{3}}+\Sigma_{33}^{\text{TDA}}(\omega)$ as functions of $\omega$ for the BH molecule. The self-energy in TDA is determined after $n$ cycles of the iterative solution of the amplitude equation, summing over the ring and ladder diagrams through the order $n+1$.
Figure 18 compares the self-energy of TDA with the exact self-energy. The former was obtained after either 20 or 21 cycles of the iterative solution of the amplitude equations, summing over ring and ladder diagrams through the order 21 or 22, respectively. As may be expected from the discussion on perturbative self-energies, TDA does not improve the overall appearance of the self-energy; rather, it degrades it.
For instance, $\bm{\Sigma}^{\text{TDA}}$ obtained after 20 cycles has qualitatively wrong (concave and convex) functional forms (outside the central bracket) because computationally it is a high-odd-order perturbation theory. In contrast, $\bm{\Sigma}^{\text{TDA}}$ after 21 cycles has qualitatively correct functional forms (as it is a high-even-order perturbation theory), but it consists of numerous vertical lines. Therefore, TDA predicts vastly different IPs and EAs depending on the number of cycles taken in the iterative solution—an artifact of calculations—and is therefore methodologically ill-defined insofar as all roots are sought. Furthermore, both of these conflicting predictions are equally meaningless. After 20 cycles, there are no real roots outside the central bracket in this graph; after 21 cycles, most roots are phantom poles with zero residues. However, this observation should not be misconstrued to mean that TDA is useless; for the principal IPs and EAs, TDA is sound and useful.
It is unsurprising for an infinite partial resummation to inherit the pathologies of higher-order perturbation theory, but this does not imply that the same pathologies plague all vertex renomalization techniques. The equation-of-motion (EOM) formalism of Green’s function theory Rowe (1968); Simons and Smith (1973); Herman et al. (1978, 1980a, 1980b) also involves an infinite partial resummation of diagrams, but it can be recast into a matrix diagonalization Baker and Pickup (1980) and may thus be free from the type of the nonconvergence problems discussed above. The random-phase approximation (RPA) Dickhoff (1999), which is a lowest-order member of the EOM approximation series, as well as the related $GW$ approximations Hedin (1965); Hybertsen and Louie (1986); Aryasetiawan and Gunnarsson (1998); Onida et al. (2002); Vlček (2019); Mejuto-Zaera and Vlček (2022) and the frequency-dependent part of the ADC self-energy Schirmer (1982); Schirmer et al. (1983) may also be more robust than TDA, although this has not been numerically confirmed nor is the convergence toward exactness implied. The Parquet method Bergli and Hjorth-Jensen (2011) sums over a related, but distinct class of diagrams and hence the present conclusion may not apply, either.
The EOM-CC method Stanton and Bartlett (1993); Stanton and Gauss (1994); Nooijen and Bartlett (1995); Bartlett et al. (1997); Stanton and Gauss (1999); Gour et al. (2005); Kamiya and Hirata (2006); Gour and Piecuch (2006); Kamiya and Hirata (2007); Hirata et al. (2000); Bartlett (2012) has been argued to be a coupled-cluster Green’s function Schönhammer and Gunnarsson (1978); Nooijen and Snijders (1992, 1993); Meissner and Bartlett (1993); Kowalski et al. (2014); Peng and Kowalski (2016); Bhaskaran-Nair et al. (2016); Peng et al. (2019, 2020); Rehr et al. (2020); Pathak et al. (2023) and can also be viewed as an infinite resummation of the propagator diagrams Hirata et al. (2015). However, this method does not suffer from the pathologies discussed above; it instead yields IPs and EAs for all (principal, satellite, and even core-ionized Vidal et al. (2019)) states that are systematically convergent at the exact values with increasing excitation level. EOM-CC separates the IP and EA sectors and its roots are obtained as eigenvalues of a non-Hermitian matrix. Like MBGF($n$), the EOM-CC roots are not guaranteed to be real, but they are almost never complex in practice.
It may be inferred that the EOM formalisms of Green’s functions, be they based on perturbation theory or coupled-cluster theory, are more robust and potentially converging because their IPs and EAs are obtained directly as eigenvalues of some matrix rather than by solving an inverse Dyson equation. In other words, the EOM formalisms may attain stability and convergence by divorcing from Green’s function theory in the Feynman–Dyson style. That the exact MBGF and MBGF(2), the latter being the only Feynman–Dyson perturbation order largely exempt from the pathologies, can also be recast into a matrix eigenvalue problem Walter et al. (1984) is consistent with this inference.
IV.2 Edge renormalization
@@
Another way of performing a diagram resummation is by renormalizing edges. By replacing all three edges in each diagram of $\delta\bm{\Sigma}^{(2)}$ by the corresponding bold-line Green’s functions $\bm{G}^{\text{Dyson}(2)}$ of Fig. 7, we include an infinite number of “row-house” diagrams appearing in the second line of Fig. 19. Furthermore, if the bold-line Green’s function (designated by $\bm{G}^{\text{sc}(2)}$) comes from this very edge-modified self-energy $\bm{\Sigma}^{\text{sc}(2)}$, as in Fig. 20, we account for another infinite set of “tower” diagrams shown in the third line of Fig. 19. This self-consistency between the self-energy and Green’s function was emphasized by Baym and Kadanoff Baym and Kadanoff (1961); Baym (1962); Kadanoff and Baym (2018); Dahlen and van Leeuwen (2005) as an essential ingredient of an approximate MBGF method that obeys conservation laws. (The first-order self-consistent MBGF method is identified as the HF theory Dickhoff (1999).) In chemistry, approximations inspired by this idea are enjoying a revival for its possible ability to describe strong correlation more accurately Surján and Szabados (1998); Holleboom and Snijders (1990); Phillips and Zgid (2014); Coveney and Tew (2023).
@@
Figure 20: The diagrammatic equation for the bold-line (renormalized) edge of the self-consistent second-order Green’s function method.
@@
This ansatz may differ somewhat from those of Van Neck et al. Van Neck et al. (1991) or of Dahlen and van Leeuwen Dahlen and van Leeuwen (2005) as it involves some additional approximations, but these differences have no impact on the following analysis. The corresponding residues are given by
@@
Figure 21 compares the self-energies of the self-consistent second-order Green’s function method after zeroth, first, and second iterative cycles. In each case, the self-energy has the qualitatively correct functional form, i.e., monotonically decreasing in each bracket separated by its singularities. This is because the self-energy expression [Eq. (32)] is isomorphic to the second-order self-energy, which has only the first-order poles. Furthermore, unlike higher-order perturbative self-energies or TDA, the brackets’ boundaries shift from one cycle to the next, and hence the satellite roots are no longer confined by brackets of the mere HF orbital energy differences.
@@
Figure 22 shows the poles and residues in the zeroth and first cycles of the self-consistent second-order Green’s function method in comparison with the exact poles and residues. The bottom panel is the same as Fig. 4, but shown in a full domain of $\omega$, and the top panel is equivalent to MBGF(2) in the diagonal approximation. In the zeroth cycle ($n=0$), the distribution of the poles agree reasonably well with that of the exact poles. The number of poles is 72 as compared with 600 exact poles including ones with zero residues. In the first cycle ($n=1$), the number of poles already reaches 4,314, far exceeding the total number (600) of ionized and electron-attached states of a FCI calculation. Furthermore, these spurious poles are not necessarily phantom mathematical roots with zero residues; they have nonzero residues and encroach on the regions where there are no exact poles. In the second cycle ($n=2$), the number of roots reaches such an astronomical value that our computer code can no longer handle, and we judged that it was not worthwhile to pursue full self-consistency. This difficulty was recognized by earlier workers, many of whom then decided to abandon this class of methods.
The bold-line self-energy diagrams in Fig. 19 is obtained by cutting one bold line (and then trimming the resulting two dangling lines) Hirata et al. (2017) of a skeleton diagram of the Luttinger–Ward functional Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Lin and Lindsey (2018) in the zero-temperature limit. Several mathematical difficulties associated with this functional have been reported recently Schäfer et al. (2013); Stan et al. (2015); Kozik et al. (2015); Rossi and Werner (2015); Schäfer et al. (2016); Tarantino et al. (2017); Gunnarsson et al. (2017).
V Analysis
V.1 Cause of nonconvergence
What is the cause of the nonconvergence? This can be answered by analyzing a model Green’s function of the form Hirata (2021b),
@@ V.1 Cause of nonconvergence
which consists of four poles. These poles are stipulated to be quadratic functions of the perturbation strength $\lambda$.
@@
The functional forms of $g(\omega)$ and $E$’s are arbitrary, but a different choice would lead to essentially the same conclusion.
@@
A truncation of this series after a finite number of terms captures the key features of the Feynman–Dyson perturbation approximations of a Green’s function or self-energy.
Figure 23: Taylor expansions up to the second order in $\lambda$ of the model Green’s function $g(\omega)$ of Eq. (37).
Figure 23 shows the zeroth-, first-, and second-order Taylor expansions of $g(\omega)$. The exact $g(\omega)$ and its zeroth-order approximation $g^{(0)}(\omega)$ have essentially the same functional forms in that they are separated into consecutive $\omega$ brackets by their singularities and within each bracket they are monotonically decreasing functions. The first-order approximation $g^{(0)}+g^{(1)}$ has a qualitatively different functional form, which is convex or concave except in the central bracket. The $g^{(0)}+g^{(1)}+g^{(2)}$ largely restores the same functional form as the exact $g(\omega)$, although the former is a rather poor approximation at many frequencies despite the fact that the second-order Taylor expansions of $E_{1}$ through $E_{4}$ are exact. These are consistent with the overall patterns of behaviors of ab initio odd- and even-order perturbative self-energies observed in the foregoing sections.
@@
Figure 24 extends this analysis to the nineteenth-order Taylor expansion. It is convergent at the exact $g(\omega)$ in some domains of $\omega$, but nonconvergent in the other domains. Generally, convergence is attained within the overlap of the bracket demarcated by the singularities of $g(\omega)$ and the one delineated by the singularities of $g^{(0)}(\omega)$. (The latter singularities are where the near-vertical rapid oscillations of the nineteenth-order Taylor expansion occur.) In this overlap, both $g(\omega)$ and $g^{(0)}(\omega)$ are in the middle “shoulder” part of the monotonically decreasing functions of $\omega$, displaying similar functional forms. Here, a Taylor expansion from $g^{(0)}(\omega)$ is convergent. Outside this overlap, either $g(\omega)$ or $g^{(0)}(\omega)$ is in the left “neck” part of the function, while the other is in the right “arm” part. They are both monotonically decreasing, but one of them has a singularity at one boundary of the domain, while the other has a singularity at the other boundary, and they have dissimilar functional forms in this sense. The $g(\omega)$ in this domain is nonanalytic and its Taylor expansion has zero radius of convergence.
@@
In a molecular Green’s function or self-energy, there are dense manifolds of singularities outside the central overlapping bracket, which encloses most principal roots and some low-energy satellite roots. Therefore, in practice, their perturbation expansions are safely convergent only in this central overlapping bracket as well as in the two terminal overlapping brackets. The central overlapping bracket is an overlap of the bracket bounded by the highest 2p1h and lowest 2h1p HF orbital energy differences and the bracket demarcated typically by the least negative and least positive singularities of the exact self-energy Sun and Bartlett (1996b). Only the roots in this central overlapping bracket, be they principal or satellite, are reliably determined by MBGF($n$).
V.2 Padé approximants
In Sec. V.1, the nonconvergence is shown to be caused by the very definition of the exact Green’s function, which is nonanalytic. Even when the positions of the poles, $E_{1},\dots,E_{4}$, are exactly expanded by second-order Taylor series, the Taylor expansion of $g(\omega)$ is not only inexact at the second order, but also nonconvergent at exactness at an infinite order. Therefore, the cause of nonconvergence is a mathematical one: The rational-function form of the exact Green’s function is unamenable to a converging Taylor expansion, even though it contains full physical information about the poles.
For this mathematical problem, the purely mathematical resummation technique of Padé approximants Bender and Orszag (1999) may prove useful. The $[M,N]$ Padé approximant of a function $f(\lambda)$ is a power-series expansion of a rational function and is defined by
$[M,N]_{f}(\lambda)=\frac{a_{0}+a_{1}\lambda+a_{2}\lambda^{2}+\dots+a_{M}\lambda^{M}}{1+b_{1}\lambda+b_{2}\lambda^{2}+\dots+b_{N}\lambda^{N}},$ (43)
whose expansion coefficients, $\{a_{0},\dots,a_{M}\}$ and $\{b_{1},\dots,b_{N}\}$, are determined such that the above agrees with the $M+N+1$-order Taylor expansion of $f(\lambda)$. Padé approximants are known to generate a rapidly converging series out of even a divergent Taylor series, often working well “beyond their proven range of applicability” Bender and Orszag (1999) at essentially no additional cost if the $M+N+1$-order Taylor expansion dominates the cost.
Figure 25: Same as Fig. 23 but for Padé approximants.
Figure 25 shows the [1,1], [1,2], and [2,2] Padé approximants of $g(\omega)$. The [2,2] Padé approximant (summing through the fifth-order Taylor series) is nearly exact and indistinguishable to the human eye from $g(\omega)$ except for a few spurious spikes. This is in contrast to Fig. 24 in which even the nineteenth-order Taylor expansion suffers from infinite errors in multiple $\omega$ domains. The remarkable performance of Padé approximants is ascribed to separately expanding the numerator and denominator of the rational-function form of $g(\omega)$.
Padé approximants were introduced to many-body perturbation energies by Goscinski Goscinski (1967) and further developed by Brändas and Goscinski Brändas and Goscinski (1970); Goscinski and Brändas (1971) and by Bartlett and Brändas Bartlett and Brändas (1972, 1973). They were applied to both convergent Laidig et al. (1985) and divergent Hirata and Bartlett (2000b) series of many-body perturbation energies, and in the latter case a rapid convergence was restored. In each case, the size-consistency is maintained Bartlett and Brändas (1972, 1973). Their applications to Green’s functions were proposed by Goscinski and Lukman Goscinski and Lukman (1970) and by Linderberg and Ratner Linderberg and Ratner (1970), but no numerical results beyond a two-site model calculation were given.
Figure 26: The third diagonal element of $\bm{\epsilon}+\bm{\Sigma}(\omega)$ as a function of $\omega$ of the BH molecule obtained by a [2,1] Padé resummation of $\bm{\Sigma}^{(n)}(\omega)$ over $1\leq n\leq 5$. The corresponding elements of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and $\bm{\epsilon}+\bm{\Sigma}^{(5)}(\omega)$ are also shown.
Figure 27: The same as Fig. 26 but for a [5,5] Padé resummation of $\bm{\Sigma}^{(n)}(\omega)$ over $1\leq n\leq 12$.
Figures 26 and 27 show the [2,1] and [5,5] Padé approximants of the perturbative self-energies of the BH molecule. They were obtained by using Eqs. (12)–(17) of Laidig and Bartlett Laidig et al. (1985) of which their “$E^{(n)}$” should be read as $\Sigma^{(n)}$ in our case. The $[N,N-1]$ and $[N,N]$ Padé approximants sum over perturbative self-energies through $12N+$ 1 and $22N+$ 2 orders, respectively, at no additional cost. The [2,1] Padé approximant (Fig. 26) not only largely eliminates the incorrect (concave and convex) shapes of $\Sigma^{(5)}$, but it also shifts the positions of the singularities in the right directions to be in better agreement with the exact self-energy. The [5,5] Padé approximant (Fig. 27) traces the exact self-energy even more closely, although spurious spikes can be seen.
[…]